Helps FIUNITI DAVUV APPILLUNS, DICONCVI Analysis, LUL SUULLIE, Taiyel IES LUDI, LUUS, LUD 4, LO5-5, LO5-6, LO5-8] Northwoo
Answer #1
Req 1
CM Ratio 40.00%
Unit Sales to break even 26400
Degree of operating leverate 3.28
CM ratio = contribution /sales
380,000/950,000
40.00%
BEP(units) = fixed cost/contribution margin per unit
264000/10
26400
Degree of operating leverage = contribution/net income
380,000/116000
3.28
Req 2 CM Ratio 28.00%
Unit Sales to break even 37714
CM ratio = contribution /sales
7./25
28.00%
BEP(units) = fixed cost/contribution margin per unit
264000/7
37714
Req 3
Number of balls 54286
BEP(units) =( fixed cost+ target income)/contribution margin per unit
(264000+116000)/7
54286
Req 4 selling price 30.00
CM ratio = 40%
selling price per unit be x
variable cost per unit is 18
so selling price should be =
40%              = (x-18)/x
.40x                  = (x -18)
x                         =18/.6
x                         = 30.00
Req 5 Selling price per unit 25
New variable cost (15*60%) 9
Contribution per unit 16
fixed expense   = 264000*200%= 528000
contribution margin ratio 64.00%
unit sales to break-even 33000 balls
Req 6A number of balls 40250 balls
(528000+116000)/16
Req6B Contribution income statement
Sales (38000*25) 950000
Variable expenses (38000*9) 342000
Contribution margin 608000
Fixed expenses 528,000
Net operating income 80,000
Degree of operating leverage 7.60
(contribution margin/net income)